OpenStudy (anonymous):

Answer the questions about the following function. f(x)=2x^2-x-1 (a) Is the point (2,25) on the graph of f? (b) If x =-2, what is f(x)? What point on the graph of f? (c) If f(x)=-1, what is x? What point(s) are on the graph of f? (d) What is the domain of f? (e) List the x-intercept(s), if any, of the graph of f. (f) List the y-intercept, if any, of the graph of f.

OpenStudy (anonymous):

did you find \(f(2)\)?

OpenStudy (anonymous):

What do you mean

OpenStudy (anonymous):

\[f(x)=2x^2-x-2\] if \((2,25)\) is on the graph, that means \(f(2)=25\) so you have to find \[f(2)=2\times 2^2-2-2\]

OpenStudy (anonymous):

if you get 25, the answer is "yes" and if you do not, the answer is "no"

OpenStudy (anonymous):

the answer was yes

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

My math is all through a website from pearson

OpenStudy (anonymous):

that may be, but \[f(2)=2\times 2^2-2-2=8-2-2=6-2=4\] and not \(25\)

OpenStudy (anonymous):

this system is a bit messed up and this is only my 3rd week in the class. I was doing good the 1st couple weeks now its getting a bit beyond what i know

OpenStudy (anonymous):

are you sure it is \[f(x)=2x^2-x-2\]?

OpenStudy (anonymous):

For part A the answer was yes

OpenStudy (anonymous):

NOw it wants to know If x=-2, what is f(x)?

OpenStudy (anonymous):

no, for part A the answer is NO

OpenStudy (anonymous):

\(f(2)=4\) not \(25\)

OpenStudy (anonymous):

it said it was right

OpenStudy (anonymous):

it isn't

OpenStudy (anonymous):

Is the point (2,5) on the graph and i marked yes and it was right

OpenStudy (anonymous):

ooh i thought you wrote \((2,25)\)

OpenStudy (anonymous):

oh no it's ok though. Now I am on part B

OpenStudy (anonymous):

\[f(x)=2x^2-x-1\] \[f(-2)=2\times (-2)^2-(-2)-1\]

OpenStudy (anonymous):

you get \[2\times 4+2-1=8+2-1=10-1=9\]

OpenStudy (anonymous):

Part B says: If x=-2, what is f(x)? f(-2)=

OpenStudy (anonymous):

right, the answer is above

OpenStudy (anonymous):

so put all of this 2×4+2−1=8+2−1=10−1=9

OpenStudy (anonymous):

no you can just put in 9

OpenStudy (anonymous):

Using this information, list a point on the graph of f. (simplify your answer. Type and ordered pair, using integers or fractions.)

OpenStudy (anonymous):

since \(f(-2)=9\) a point on the graph is \((-2,9)\)

OpenStudy (anonymous):

Part C If f(x)=-1, what is x?

OpenStudy (anonymous):

you have to solve \[2x^2-x-1=-1\] do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

add one to both sides and get \[2x^2-x=0\] then factor

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

also \(0\) there are two solutions

OpenStudy (anonymous):

so (1/2,0)

OpenStudy (anonymous):

the two solutions are \(x=0\) and \(x=\frac{1}{2}\)

OpenStudy (anonymous):

is that how i would type it in

OpenStudy (anonymous):

it is not an ordered pair, i have no idea what pearson wants you to write

OpenStudy (anonymous):

ahh the 'points on the graph" are \((0,-1)\) and \((\frac{1}{2},-1)\)

OpenStudy (anonymous):

ok what is the domain of f

OpenStudy (anonymous):

type it interval notation

OpenStudy (anonymous):

I can't figure the domain out